OpenStudy (chillhill):

factorise: 3x^2-6x-9=0. I got the answer as (x-3)(x+9) but someone else got 3(x-3)(x+1), Which one is correct?

OpenStudy (will.h):

what is the GCF here?

OpenStudy (will.h):

(Greatest common factor|)

OpenStudy (agent0smith):

You can pretty easily check your work here... expand this out, (x-3)(x+9), and do you get back the original?

OpenStudy (will.h):

3x^2 - 6x - 9 has a GCF of 3 divide each term by 3 3x^2/3 = x^2 -6x/3 = -2x -9/3 = -3 so we now have 3(x^2 - 2x - 3) now can we factor any further? yes! factoring by groups 2 numbers when multiplied must equal to -2 and their sum is equal to -2 that would be 1 and -3 3(x+1) (x-3 )

OpenStudy (will.h):

when multiplied must equal to -3 not -2 sorry

OpenStudy (austin.l):

I wish you could give out multiple medals. Agent's method is the quickest way to check and Will has the maths down. Both are good methods for future reference.